how many ml of water at 10 degrees celsius (2 significant figures) must be added to 75 ml of water at 35 degrees celsius to obtain a final temperature of 19 degrees celsius?


1
Expert's answer
2021-09-02T02:01:38-0400

Q=mCΔT

; where

m is the mass of the substance,

C is the specific heat of the substance, and

ΔT is the difference in temperature.

35 -19 = 16

19-10 = 9

16 + 9 = 25

0.075 × 25 × 4.189 = 7.85 ml